### v2

# Topological Stability and Latschev-type Reconstruction Theorems for $\boldsymbol{\mathrm{CAT}(\kappa)}$ Spaces

Abstract.
We consider the problem of homotopy-type reconstruction of compact shapes $X\subset\mathbb{R}^{N}$ that are $\mathrm{CAT}(\kappa)$ in the intrinsic length metric. The reconstructed spaces are in the form of Vietoris–Rips complexes computed from a compact sample $S$, Hausdorff–close to the unknown shape $X$. Instead of the Euclidean metric on the sample, our reconstruction technique leverages a path-based metric to compute these complexes. As naturally emerging in the framework of reconstruction, we also study the Gromov–Hausdorff topological stability and finiteness problem for general compact $\mathrm{CAT}(\kappa)$ spaces. Our techniques provide novel sampling conditions as an alternative to the existing and commonly used techniques using weak feature size and $\mu$–reach. In particular, we introduce a new parameter, called the restricted distortion, which is a generalization of the well-known global distortion of embedding. We show examples of Euclidean subspaces, for which the known parameters such as the reach, $\mu$–reach and weak features size vanish, whereas the restricted distortion is finite, making our reconstruction results applicable for such spaces.
Key words and phrases:
Vietoris–Rips complex, shape reconstruction, $\mathrm{CAT}(\kappa)$ spaces, Latschev’s theorem, homotopy equivalence
2020 Mathematics Subject Classification:
55P10 (Primary), 55N31, 54E35 (Secondary)
1. Introduction
In the last two decades, the problem of shape reconstruction has received increasing attention in theoretical and applied communities alike. Although the shape reconstruction problem comes in many forms, our study focuses on the reconstruction of a homotopy-type for both abstract and Euclidean embedded shapes under the Gromov–Hausdorff $d_{GH}$ and Hausdorff distance $d_{H}$, respectively. The former type models the hidden underlying shape as an abstract metric space $(X,d_{X})$ and a sample $(S,d_{S})$ as a (possibly finite) metric space having a small Gromov–Hausdorff (Definition 2.5) to $X$. Whereas, the latter type assumes the unknown shape to be a Euclidean subset $X\subset\mathbb{R}^{N}$ and the sample a Euclidean subset $S\subset\mathbb{R}^{N}$ close to $X$ in the Hausdorff distance (Definition 2.4). In both cases, the goal is to produce (from the sample $S$) a topological space $\widetilde{X}$—called the reconstruction of the underlying shape $X$—such that $\widetilde{X}$ is equivalent to $X$ in a given topological sense, i.e. typically up to a homeomorphism $\cong$, diffeomorphism111in the case $X$ is a differentiable manifold, or a homotopy equivalence $\simeq$, where the third is the weakest among all of them222diffeomorphism $\Rightarrow$ homeomorphism $\Rightarrow$ homotopy equivalence.
If the underlying space $X$ is a differentiable manifold333realized via a specific embedding into $\mathbb{R}^{N}$, there is a large body of literature on the reconstruction problem and well-developed algorithms for the extraction of $\widetilde{X}$. They are commonly referred to as manifold learning techniques as they find direct applications to machine learning, more widely known as AI. In this brief introduction, we only point to a fraction of existing literature in [49, 23, 24, 25, 28, 9, 26, 54], where the authors consider the problem of $C^{2}$–manifold reconstruction based on a finite, possibly noisy sample. The key property of $n$–manifolds444[49, 23, 24, 28, 9] focus primarily on closed $n$–manifolds which allows for their effective reconstruction [23, 24, 28] is the fact that they are defined via a differentiable structure: a set of smooth charts (or local parametrizations whose domains are open subsets of $\mathbb{R}^{n}$) known as atlas. Equivalently, Euclidean submanifolds can be obtained as regular level sets of $\mathbb{R}^{m}$–valued functions. Such a convenient description is not available for more general shapes that manifest irregular features like sharp corners, branches, cusps, and multi-dimensional stratification. Examples include embedded graphs555for example road networks (or more generally simplicial complexes), pinched manifolds, manifolds with boundary, stratified manifolds with strata of various dimensions, or more general compact Euclidean subsets. We also note that closed submanifolds of $\mathbb{R}^{N}$ have a positive reach $r(X)$ [22], which is a critical property required by the existing manifold reconstruction schemes [23, 24, 28]. However, the irregular shapes mentioned above, typically do not meet this assumption. The class of irregular shapes considered in the current work is known, in metric geometry [29, 12, 11], as $\mathrm{CAT}(\kappa)$ spaces (see Section 2.2). For such irregular shapes, lacking a differentiable structure, one can only hope to obtain the reconstruction $\widetilde{X}$ as a geometric realization of an abstract simplicial complex, homeomorphic or homotopy equivalent to $X$.
Each reconstruction technique relies on a set of sampling parameters, which usually depends on the geometry and topology of the hidden shape $X$. Some of the above mentioned reconstruction works use reach $r(X)$ [24, 25, 28, 43, 15, 40] and for more general shapes, the $\mu$–reach $r_{\mu}(X)$ [14, 6, 33]; see Section 6 for a comparison. Alternatively, other works [21, 39] use convexity radius $\rho(X)$ (Definitions 3.1) and global distortion $\delta(X)$ of the embedding (Definition 4.2). In this paper, we generalize the already existing notion of global distortion $\delta(X)$ (Definition 4.2) [29] to introduce a novel sampling parameter: restricted distortion $\delta^{\varepsilon}_{R}(X)$ for $\varepsilon,R>0$. The primary advantage of the new parameter $\delta^{\varepsilon}_{R}(X)$ is that it attains a finite value for various shapes with vanishing $\mu$–reach (Section 6). Our main Euclidean shape reconstruction theorem is the following.
Theorem A (Reconstruction of $\mathrm{CAT}(\kappa)$ Subspaces of $\mathbb{R}^{N}$).
Let $X$ be a compact $\mathrm{CAT}(\kappa)$ subspace of $\mathbb{R}^{N}$ and let $\xi\in\left(0,\tfrac{1}{14}\right]$ be a fixed number. For any $\beta$ satisfying $0<\beta<\frac{1}{1+2\xi}\Delta(X)$, if we choose $0<\varepsilon\leq\beta$ such that $\delta^{\varepsilon}_{2\xi\beta}(X)\leq 1+\left(\frac{\xi}{1+\xi}\right)$, then for any compact subset $S\subset\mathbb{R}^{N}$
$d_{H}(X,S)<\tfrac{1}{2}\xi\varepsilon\quad\Rightarrow\quad\mathcal{R}^{\varepsilon}_{\beta}(S)\simeq X.$
The reconstructed shape, denoted as $\widetilde{X}=\mathcal{R}^{\varepsilon}_{\beta}(S)$, presented in the theorem above, corresponds to the geometric representation of the Vietoris–Rips complex associated with a metric space $(S,d^{\varepsilon}_{S})$. Here, $d^{\varepsilon}_{S}$ is defined as a path metric (see Definition 4.4) for the sample metric space $(S,d^{\varepsilon}_{S})$. The parameter $\Delta(X)$ incorporates both the convexity radius $\rho(X)$ and the curvature limit $\kappa$ of the space $X$, as specified in (2). The adoption of the path metric is essential because traditional Euclidean Vietoris–Rips (and Čech) complexes often do not preserve topological properties, a limitation highlighted in works such as [21].
Based on the inherent geometry of the shape $X$, Theorem A specifies a range of values for the parameter $\beta$ that ensures the homotopy equivalence of these complexes to $X$. The demonstration of Theorem A is twofold. Initially, it expands upon the findings from [40] to include $\mathrm{CAT}(\kappa)$ spaces, thereby broadening the applicability of the well-established Latschev’s Theorem, referenced in [36].
Theorem B (Quantitative Latschev’s Theorem for Compact $\mathrm{CAT}(\kappa)$ Spaces).
Let $(X,d_{X})$ be a compact $\mathrm{CAT}(\kappa)$ space and $\xi\in\left(0,\tfrac{1}{14}\right]$ a fixed number. For any $\beta$ satisfying $0<\beta<\frac{\Delta(X)}{(1+2\xi)}$, if a compact metric space $(S,d_{S})$ is a $(\xi\beta,\beta)$-approximation of $X$, then $\mathcal{R}_{\beta}(S)\simeq X$. In particular,
$d_{GH}(X,S)<\xi\beta\quad\Rightarrow\quad\mathcal{R}_{\beta}(S)\simeq X,$
where $\mathcal{R}_{\beta}(S)$ is the geometric realization of the Vietoris–Rips complex of $(S,d_{S})$ at scale $\beta$.
The proof of Theorem B is presented in Section 3.3 and is structurally similar to its predecessor (Theorem 1.1) proven in [40] since it relies on the well known Whitehead Theorem [31]. A discernible difference is that the weaker666than the Gromov–Hausdorff closeness notion of the $(\varepsilon,R)$–approximation (Definition 2.6) is employed in Theorem B. While this theorem provides foundational insights, it alone does not suffice for the comprehensive reconstruction results aimed for in Theorem A, which stands as the centerpiece of this paper. Nevertheless, Theorem B facilitates the exploration of topological stability within $\mathrm{CAT}(\kappa)$ spaces.
It should be noted that Gromov–Hausdorff limits are generally not topologically faithful, as outlined in sources like [29, 12]. However, stability results have been established within specific classes of spaces. One prominent example is Gromov’s Theorem [29, Theorem 8.19], which guarantees stability (up to a diffeomorphism) for complete Riemannian $n$–manifolds that have two-sided sectional curvature bounds and a uniform minimum on their injectivity radii.
Additionally, Alexandrov spaces characterized by a lower bound on curvature, fixed dimensions, and a diameter constraint also exhibit stability, in this case up to a homeomorphism, as demonstrated in studies like [13, 44] and for manifolds in [30]. Thanks to Theorem B, here we obtain the following homotopy-type stability result for $\mathrm{CAT}(\kappa)$ spaces.
Theorem C (Topological Stability of compact $\mathrm{CAT}(\kappa)$ Spaces).
Let $X,X^{\prime}$ be compact $\mathrm{CAT}(\kappa)$ spaces. For any $\xi\in\left(0,\tfrac{1}{14}\right]$, let $\beta$ be a number satisfying $0<\beta<\frac{1}{(1+2\xi)}\min\{\Delta(X)$, $\ \Delta(X^{\prime})\}$. If $X$ and $X^{\prime}$ are $(\xi\beta,\beta)$-approximations of each other, then $X$ and $X^{\prime}$ are homotopy equivalent. In particular,
$d_{GH}(X,X^{\prime})<\xi\beta\quad\Rightarrow\quad X\simeq X^{\prime}.$
Homotopy-type stability in the setting of Riemannian manifolds and Finsler manifolds has previously been obtained in [53] and [55]. For general $\operatorname{LGC}$–spaces (i.e. locally geometrically contractible spaces) with bounded topological dimension, homotopy-type stability was obtained in [46]. The $\mathrm{CAT}(\kappa)$ spaces, considered here, are $\operatorname{LGC}$–spaces (due to the lower bound for the convexity radius), however in contrast to results in [46], Theorem C does not impose restrictions on the topological dimension of spaces.
A typical corollary of a stability result of this kind is a quantitative analog of the well-known Cheeger’s finiteness theorem for Riemannian manifolds [16] (see also [45, grove1990geometricfinitness, 46]). In particular, Theorem C implies
Corollary D.
In a class of compact $\mathrm{CAT}(\kappa)$ spaces, with a fixed upper curvature bound $\kappa$, and a uniform lower bound for convexity radius, any precompact family admits only finitely many homotopy types.
The second component in the proof of Theorem A is the following approximation result for compact geodesic subspaces of $\mathbb{R}^{N}$ by a Hausdorff-close Euclidean sample.
Theorem E ($(\varepsilon,R)$–approximation Theorem).
Let $X$ be a geodesic subspace of $\mathbb{R}^{N}$. Let $\xi\in\left(0,1\right)$ and $\beta>0$ be fixed numbers. If we choose777it is possible to choose such a small $\varepsilon$, thanks to Proposition 5.3. $0<\varepsilon\leq\beta$ such that $\delta^{\varepsilon}_{2\xi\beta}(X)\leq 1+\left(\frac{\xi}{1+\xi}\right)$, then for any compact subset $S\subset\mathbb{R}^{N}$ with $d_{H}(X,S)<\frac{1}{2}\xi\varepsilon$, the metric space $(S,d^{\varepsilon}_{S})$ is an $(\xi\beta,\beta)$-approximation of $(X,d^{L}_{X})$.
A key element of Theorem E is the employment of the path metric $d^{\varepsilon}_{S}$ of the sample $S$. The notion of the path metric in this context was previously introduced in [21, 39]. In this study, we further our understanding of these metrics. In particular, we show in Theorem 4.8 that they are stable under Hausdorff perturbation. Moreover in Theorem 4.9, the path metrics on a compact geodesic space $X$ have been shown to converge to the intrinsic metric of $X$.
1.1. Background on Latschev’s Theorem
Hausmann proved in [32] that a closed Riemannian manifold is homotopy equivalent to its own Vietoris–Rips complex for a sufficiently small scale (see Theorem 3.2). The result quite naturally elicits the question of finite reconstruction: Under what conditions is a Riemannian manifold homotopy equivalent to the Vietoris–Rips complex of a finite but dense sample? In [36], Latschev later provided an answer to the above manifold reconstruction problem satisfactorily, and also in a little more general context of density measured in the Gromov–Hausdorff distance (Definition 2.5). Latschev’s theorem guarantees: For a closed Riemannian manifold $X$, there exists a positive number $\epsilon_{0}$ such that for any scale $0<\beta\leq\epsilon_{0}$ there exists some $\delta>0$ such that for every metric space $S$ with Gromov–Hausdorff distance to $X$ less than $\delta$, the Vietoris–Rips complex $\mathcal{R}_{\beta}(S)$ is homotopy equivalent to $X$. Although the result is purely qualitative and existential, its relevance can be ascribed to the emphasis that the quantity $\epsilon_{0}$ must depend solely on the intrinsic geometry of $X$. Recently, the new quantitative version of Latschev’s theorem has been proved by Majhi [40].
Theorem 1.1 (Quantitative Latschev’s Theorem for Riemannian manifolds [40]).
Let $X$ be a closed, connected Riemannian manifold. Let $S$ be a compact metric space and $\beta>0$ a number such that
$\frac{1}{\xi}d_{GH}(X,S)<\beta<\frac{\Delta(X)}{1+2\xi}$
for some $\xi\in\left(0,\tfrac{1}{14}\right]$. Then, $\mathcal{R}_{\beta}(S)\simeq X$.
The result, indeed, demystifies the constants in Latschev’s theorem:
$\epsilon_{0}=\frac{\Delta(X)}{1+2\xi}\text{ and }\delta=\xi\beta.$
In the same spirit, a Latschev-type result was also presented in [40] for the reconstruction of Euclidean submanifolds from a Hausdorff-close sample. The reach $r(X)$ of the manifold $X$ has been used as the sampling parameter; see [22] for definition.
Theorem 1.2 (Submanifold Reconstruction under Hausdorff Distance [40]).
Let $X\subset\mathbb{R}^{d}$ be a closed, connected Euclidean submanifold. Let $S\subset\mathbb{R}^{d}$ be a compact subset and $\beta>0$ a number such that
$\frac{1}{\xi}d_{H}(X,S)<\beta\leq\frac{3(1+2\xi)(1-14\xi)}{8(1-2\xi)^{2}}r(X)$
for some $\xi\in\left(0,\tfrac{1}{14}\right)$. Then, $\mathcal{R}_{\beta}(S)\simeq X$. Here, $\mathcal{R}_{\beta}(S)$ denotes the Vietoris–Rips complex of $S$ under the Euclidean distance.
1.2. Related works on the reconstruction of homotopy type
The homotopy-type reconstruction problem has been investigated for various shape classes, we list a fraction of the existing literature in [43, 14, 15, 6]. In the seminal work of Niyogi et al. [43] the authors showed for a sufficiently small radius, controlled by the reach $r(X)$, the Euclidean thickening of a dense enough sample deformation retracts onto the submanifold $X\subset\mathbb{R}^{N}$. Beyond smooth submanifolds, a broader class of shapes includes the compact Euclidean subsets with a positive reach [22], more generally positive $\mu$-reach $r_{\mu}(X)$ [14] for an appropriately chosen value of the parameter $\mu$ (see Section 6). The known reconstruction results for such shapes were obtained in [6] and [33] using both Vietoris–Rips and Čech complexes. Another very important class of shapes is embedded metric graphs see the work in [39, 37] for reconstruction via the Vietoris–Rips complexes.
1.3. Acknowledgements
The first and third author acknowledge the partial support by Louisiana Board of Regents Targeted Enhancement Grant 090ENH-21.
2. Preliminaries
In this section, we provide some preliminary notation and definitions used throughout the paper.
2.1. Length Spaces
Let $(X,d)$ be a metric space. For a (continuous) path $\gamma:[a,b]\to X$ in $X$, we first define the notion of its length, denoted $L_{d}(\gamma)$, induced by the metric $d$.
We call $T=\{t_{i}\}_{i=0}^{k}$ a partition of $[a,b]$ if:
$a=t_{0}<t_{1}<t_{2}<\ldots<t_{k}=b.$
Then, the length of $\gamma$ can be defined as:
$L_{d}(\gamma)\coloneqq\sup\sum_{i=0}^{k-1}d(\gamma(t_{i}),\gamma(t_{i+1})),$
where the supremum is taken over all partitions $T$ of $[a,b]$. A path $\gamma$ is called rectifiable if its length is finite. When the choice of the metric $d$ is obvious from the context, we drop the subscript $d$ and denote the induced length simply by $L(\cdot)$. We note from [12, Proposition 2.3.4] that $L$ is lower semi-continuous, a property that finds its use later in the paper.
Lemma 2.1 (Lower Semi-Continuity).
Let $(X,d)$ be a metric space. If a sequence of rectifiable paths $\{\gamma_{i}\colon[a,b]\to X\}_{i=1}^{\infty}$ and a path $\gamma:[a,b]\to X$ are such that $\lim\limits_{i\to\infty}d(\gamma_{i}(t),\gamma(t))=0$ for every $t\in[a,b]$, then lim $\liminf L(\gamma_{i})\geq L(\gamma)$.
The notion of the length of paths in $X$, in turn, associates a distance between any two points of $X$.
Definition 2.2 (Induced Length Metric).
For any two points $x_{1},x_{2}\in X$, the induced length metric, denoted $d^{L}(x_{1},x_{2})$, defined by the infimum of lengths of paths joining $x_{1},x_{2}$:
$d^{L}(x_{1},x_{2})\coloneqq\inf\left\{L(\gamma)\mid\gamma:[a,b]\to X,\gamma(a)=x_{1},\gamma(b)=x_{2}\right\}.$
Although not always finite, $d^{L}$ defines a metric on $X$. A metric space $(X,d)$ is called a length space and its metric $d$ is length metric if $d$ coincides with the induced length metric $d^{L}$. For any metric space $(X,d)$, its induced length metric space $(X,d^{L})$ is always a length space. Moreover, $(X,d)$ is called a complete length space if for any two points $x_{1},x_{2}\in X$, there exists a shortest path $\gamma:[a,b]\to X$ joining them, i.e., the induced distance $d^{L}(x_{1},x_{2})=L(\gamma)$.
A subset $A$ of a length space $(X,d)$ is called convex if for any two points $a_{1},a_{2}\in A$, there exists a unique shortest path joining $a_{1}$ and $a_{2}$ lies entirely in $A$.
2.2. $\boldsymbol{\mathrm{CAT}(\kappa)}$ Spaces
Alexandrov originally studied these spaces in the early 1900s to generalize the notion of Riemannian curvature to length spaces [4, 5]. For $\kappa\in\mathbb{R}$, a $\mathrm{CAT}(\kappa)$ space $X$ is a complete length space such that the geodesic triangles in $X$ are not fatter than corresponding comparison triangles in the model surface of constant sectional curvature $\kappa$. Consequently, the class includes closed Riemannian manifolds, locally-finite metric graphs and various stratified spaces. The following formal definition is in the style of [12, 35]. The only difference is we always assume that a $\mathrm{CAT}(\kappa)$ is a complete length space.
Definition 2.3 ($\mathrm{CAT}(\kappa)$ Space).
For $\kappa\in\mathbb{R}$, a complete length space $(X,d)$ is called a $\mathrm{CAT}(\kappa)$ space if for any $x_{1},x_{2},x_{3}\in X$ (in case $\kappa>0$, additionally $d(x_{1},x_{2})+d(x_{2},x_{3})+d(x_{3},x_{1})<2\pi/\sqrt{\kappa}$ ) the following triangle comparison property holds: $d(y,x_{2})\leq d(\overline{y},\overline{x}_{2})$ for any point $y$ on a shortest path joining $x_{1}$ and $x_{3}$, where $\overline{x}_{i}$’s are points on the surface of constant curvature $\kappa$ with $d(\overline{x}_{i},\overline{x}_{j})=d(x_{i},x_{j})$ and $\overline{y}$ is the point on the shortest path joining $\overline{x}_{1}$ and $\overline{x}_{3}$ with $d(\overline{x}_{1},\overline{y})=d(x_{1},y)$.
2.3. Hausdorff and Gromov–Hausdorff Distances
Let $A$ and $B$ be compact, non-empty subsets of a metric space $(X,d)$. Roughly speaking, the Hausdorff distance between $A$ and $B$ is the infimum radius $\eta>0$ that suffices for the thickenings $A^{\eta}$ and $B^{\eta}$ to cover $B$ and $A$, respectively. Here, the thickening are taken in $X$ under the metric $d$, e.g,
$A^{\eta}\coloneqq\bigcup_{a\in A}\{x\in X\mid d(x,a)<\eta\}.$
More concretely, we have the following definition.
Definition 2.4 (Hausdorff Distance).
The Hausdorff distance between the $A$ and $B$, denoted $d^{X}_{H}(A,B)$, is defined as
$d^{X}_{H}(A,B)\coloneqq\inf\{\eta>0\mid B\subset A^{\eta}\text{ and }A\subset B^{\eta}\}.$
In case $X\subset\mathbb{R}^{N}$ and $A,B,X$ are all equipped with the Euclidean metric, we simply write $d_{H}(A,B)$.
A correspondence $\mathscr{C}$ between any two metric spaces $(X,d_{X})$ and $(Y,d_{Y})$ is defined as a subset of $X\times Y$ such that
  1. (a)
    for any $x\in X$, there exists $y\in Y$ such that $(x,y)\in\mathscr{C}$, and
  2. (b)
    for any $y\in Y$, there exists $x\in X$ such that $(x,y)\in\mathscr{C}$.
We denote the set of all correspondences between $X,Y$ by $\mathscr{C}(X,Y)$. The distortion of a correspondence $\mathscr{C}\in\mathscr{C}(X,Y)$ is defined as:
$\mathrm{dist}(\mathscr{C})\coloneqq\sup_{(x_{1},y_{1}),(x_{2},y_{2})\in\mathscr{C}}|d_{X}(x_{1},x_{2})-d_{Y}(y_{1},y_{2})|.$
Definition 2.5 (Gromov–Hausdorff Distance).
Let $(X,d_{X})$ and $(Y,d_{Y})$ be two compact metric spaces. The Gromov–Hausdorff distance between $X$ and $Y$, denoted by $d_{GH}(X,Y)$, is defined as:
$d_{GH}(X,Y)\coloneqq\frac{1}{2}\left[\inf_{\mathscr{C}\in\mathscr{C}(X,Y)}\mathrm{dist}(\mathscr{C})\right].$
A slightly weaker and more local notion is $(\varepsilon,R)$-approximation for constants $\varepsilon,R>0$. The notion has already been used in [1] in the context of reconstruction. However, we use a slightly different version to match the traditional definition of the Gromov–Hausdorff distance.
Definition 2.6 ($(\varepsilon,R)$-Approximation).
For $\varepsilon,R$ positive, two metric spaces $(X,d_{X})$ and $(Y,d_{Y})$ are said to be an $(\varepsilon,R)$-approximation of each other if there exists a correspondence $\mathscr{C}\subset X\times Y$ such that for any $(x_{1},y_{1}),(x_{2},y_{2})\in\mathscr{C}$ with $\min\{d_{X}(x_{1},x_{2}),d_{Y}(y_{1},y_{2})\}\leq R$, we have
$|d_{X}(x_{1},x_{2})-d_{Y}(y_{1},y_{2})|\leq 2\varepsilon.$
To see that this is a weaker notion, we note that $d_{GH}(X,Y)<\varepsilon$ implies that the two metric spaces are $(\varepsilon,R)$ for any $R>0$. However, an $(\varepsilon,R)$-approximation may not always imply $\varepsilon$-closeness in the Gromov–Hausdorff distance; see the following example.
Example 2.7.
[…]
2.4. Abstract Simplicial Complexes and Geometric Realizations
Definition 2.8 (Abstract Simplicial Complex).
An abstract simplicial complex is a set $\mathscr{K}$ closed under subset inclusion. That is, for each non-empty, finite $\sigma\in\mathscr{K}$, if $\sigma^{\prime}\subset\sigma$, then $\sigma^{\prime}\in\mathscr{K}$.
Each $\sigma\in\mathscr{K}$ is called a simplex of $\mathscr{K}$; a subset $\sigma^{\prime}\subset\sigma$ is called a face of the $\sigma$. If the cardinality of a simplex $\sigma\in\mathscr{K}$ is $k+1$, then we call it a $k$-simplex and define its dimension, denoted $\dim(\sigma)$, to be $k$. The collection of all 00-dimensional simplices is called the vertex set of $\mathscr{K}$.
Note $\mathscr{K}$ may have sufficient combinatorial structure for our purposes, but not enough topological structure. We may add such topological structure by defining a geometric realization of $\mathscr{K}$, denoted as $|\mathscr{K}|$ to be the topological space $\mathscr{K}$ under some choice of a topology on $\mathscr{K}$; find more details in [48]. For notational convenience, we denote by $\mathscr{K}$ both an abstract simplicial complex and its geometric realization.
A simplicial complex $\mathscr{K}$ is called a pure $m$–complex if every simplex of $\mathscr{K}$ is a face of an $m$–simplex. A simplicial complex $\mathscr{K}$ is called a flag complex if $\sigma$ is a simplex of $\mathscr{K}$ whenever every pair of vertices in $\sigma$ is a simplex of $\mathscr{K}$.
Definition 2.9 (Vietoris–Rips Complex).
Let $(X,d)$ be a metric space. Then, the Vietoris–Rips complex of $X$ at scale $\beta>0$, denoted $\mathcal{R}_{\beta}(X)$, is an abstract simplicial complex such that the $n$-simplex $[x_{0},x_{1},\ldots,x_{n}]$ is in $\mathcal{R}_{\beta}(X)$ if and only if $d(x_{i},x_{j})<\beta$ for all $0\leq i,j\leq n$.
2.5. Barycentric Subdivision
The barycenter, denoted $\widehat{\sigma}_{m}$, of an $m$-simplex $\sigma_{m}=[v_{0}$, $v_{1}$, $\ldots$, $v_{m}]$ of $\mathscr{K}$ is the point of $\langle\sigma_{m}\rangle$ such that $\widehat{\sigma}_{m}(v_{i})=\frac{1}{m+1}$ for all $0\leq i\leq m$. Using linearity of simplices, a more convenient way of writing this is:
$\widehat{\sigma}_{m}=\sum_{i=0}^{m}\frac{1}{m+1}v_{i}.$ (1)
For a simplicial complex $\mathscr{K}$, its barycentric subdivision, denoted $\operatorname{sd}{\mathscr{K}}$, is a simplicial complex such that
  1. (i)
    the vertices of $\operatorname{sd}{\mathscr{K}}$ are the barycenters of simplices of $\mathscr{K}$
  2. (ii)
    the simplices of $\operatorname{sd}{\mathscr{K}}$ are (non-empty) finite sets $[\widehat{\sigma}_{0},\widehat{\sigma}_{1},\ldots,\widehat{\sigma}_{m}]$ such that $\sigma_{i-1}\prec\sigma_{i}$ for $1\leq i\leq m$ and $\sigma_{i}\in\mathscr{K}$
  3. (iii)
    the linear map $\operatorname{sd}\colon\operatorname{sd}{\mathscr{K}}\to\mathscr{K}$ sending each vertex of $\operatorname{sd}{\mathscr{K}}$ to the corresponding point of $\mathscr{K}$ is a homeomorphism.
The following result from [39] proves a crucial ingredient for our proofs.
Lemma 2.10 (Commuting Diagram [39]).
Let $\mathscr{K}$ be a pure $m$–complex and $\mathscr{L}$ a flag complex. Let $f\colon\mathscr{K}\to\mathscr{L}$ and $g\colon\operatorname{sd}\mathscr{K}\to\mathscr{L}$ be simplicial maps such that
  1. (1)
    $g(v)=f(v)$ for every vertex $v$ of $\mathscr{K}$,
  2. (2)
    $f(\sigma)\cup g(\widehat{\sigma})$ is a simplex of $\mathscr{L}$ whenever $\sigma$ is a simplex of $\mathscr{K}$.
Then, the following diagram commutes up to homotopy:
$\operatorname{sd}{\mathscr{K}}$$\mathscr{K}$$\mathscr{L}$${\operatorname{sd}^{-1}}$$g$$f$
where $\operatorname{sd}$ is the linear homeomorphism sending each vertex of $\operatorname{sd}{\mathscr{K}}$ to the corresponding point of $\mathscr{K}$.
Theorem 2.11 (Whitehead Theorem, [31]).
Let $X$ and $Y$ be topological spaces, homotopy equivalent to CW–complexes888which are allowed to be uncountably infinite, as we encounter e.g. in Theorem 3.2 and in Theorem A.. If $f:X\to Y$ is a map inducing isomorphisms on all homotopy groups, i.e., $\pi_{m}(f_{*}):\pi_{m}(X)\to\pi_{m}(Y)$ is an isomorphism for all $m\geq 0$, then $f$ is a homotopy equivalence.
3. Latschev’s Theorem For $\mathrm{CAT}(\kappa)$ Spaces
This section extends Latschev’s theorem to $\mathrm{CAT}(\kappa)$ spaces (Definition 2.3). In addition, we discuss how notions like convexity radius and results like Jung’s and Hausmann’s theorem make sense for such spaces.
3.1. Hausmann’s Theorem
We begin with the definition of the convexity radius of a length space.
Definition 3.1 (Convexity Radius).
The convexity radius of a length space $(X,d)$, denoted $\rho(X)$, is defined as the infimum of the set of radii of the largest convex metric balls across the points of $X$. Formally,
$\rho(X)=\inf_{x\in X}\sup\;\{r\geq 0\mid\mathbb{B}_{X}(x,s)\text{ is convex for all }0<s<r\}.$
Due to comparison triangles in small neighborhoods, a metric ball in a $\mathrm{CAT}(\kappa)$ space is always convex for a sufficiently small radius ([7, Lemma 3.4]). For a compact $\mathrm{CAT}(\kappa)$ space, as a consequence, its convexity radius $\rho(X)>0$.
Although Hausmann defines of convexity radius in the context of Riemannian manifolds, a remark states that the definition and the results in [32] are equally valid for length spaces. We also remark that Hausmann’s definition is slightly different from ours. However, the alteration does not tamper with the veracity of Hausmann’s theorem. We now state Hausmann’s theorem for $\mathrm{CAT}(\kappa)$ spaces.
Theorem 3.2 (Hausmann’s Theorem [32]).
Let $X$ be a compact $\mathrm{CAT}(\kappa)$ space with convexity radius $\rho(X)$. For any $0<\beta<\rho(X)$, we have a homotopy equivalence $\mathcal{R}_{\beta}(X)\simeq X$.
Furthermore, if $0<\beta\leq\beta^{\prime}<\rho(X)$, then the inclusion $\mathcal{R}_{\beta}(X)\hookrightarrow\mathcal{R}_{\beta^{\prime}}(X)$ is a homotopy equivalence [51].
In this context, we note the works of [2, 3], [52], and [38] for variations and extensions of Hausmann’s theorem and alternative proofs.
Throughout the section, we assume that $(X,d)$ is a compact $\mathrm{CAT}(\kappa)$ space; consequently, $\rho(X)>0$. We define the quantity
$\Delta(X)\coloneqq\begin{cases}\min\left\{\frac{\pi}{4\sqrt{\kappa}},\rho(X)\right\},&\text{ if }\kappa>0,\\ \rho(X),&\text{ if }\kappa\leq 0.\end{cases}$ (2)
3.2. Jung’s Theorem
We first state the classical Jung’s Theorem in the Euclidean space, then in $\mathrm{CAT}(\kappa)$ spaces. For a bounded subset $A$ of any metric space $(X,d)$, we define its circumradius to be the radius of the smallest (closed) metric ball enclosing $A$. More formally,
$\operatorname{rad}(A)\coloneqq\inf_{x\in X}\ \sup_{a\in A}d(a,x).$
A point $x\in X$ satisfying $\max_{a\in A}d(a,x)=\operatorname{rad}(A)$ is called a circumcenter of $A$, and is denoted by $c(A)$. For $A$ compact, a circumcenter may not uniquely exist. We warn the reader that these terms may not always match their usual meaning in plane geometry.
For a subset $A\subset\mathbb{R}^{N}$ with exactly $n+1$ distinct points, the classical Jung’s theorem [17, Theorem 2.6] states that $c(A)$ is unique and $\operatorname{rad}(A)\leq\sqrt{\frac{n}{2(n+1)}}\operatorname{diam}(A)$.
The above bound on the circumradius has been extended in both directions; see for example [19, 20, 18, 35]. Particularly relevant is the extension of Jung’s theorem to $\mathrm{CAT}(\kappa)$ spaces [35].
Theorem 3.3 (Generalized Jung’s Theorem [35]).
Let $X$ be a complete $\mathrm{CAT}(\kappa)$ space with a non-empty subset $A$ of $X$ with at most $n+1$ distinct points for any positive integer $n$. In the case where $\kappa>0$, assume the circumradius of $A$, $\operatorname{rad}(A)<\frac{\pi}{2\sqrt{\kappa}}$. Then, there is a unique circumcenter $c(A)$ in $X$. Moreover,
$\operatorname{diam}(A)\geq\begin{cases}\frac{2}{\sqrt{\kappa}}\sin^{-1}\left(\sqrt{\frac{n+1}{2n}}\sin{(\sqrt{\kappa}\operatorname{rad}(A))}\right)&\text{ if }\kappa>0,\\ 2\sqrt{\dfrac{n+1}{2n}}\operatorname{rad}(A)&\text{ if }\kappa=0,\\ \frac{2}{\sqrt{-\kappa}}\sinh^{-1}\left(\sqrt{\frac{n+1}{2n}}\sinh{(\sqrt{-\kappa}\operatorname{rad}(A))}\right)&\text{ if }\kappa<0.\end{cases}$ (3)
The following lemma provides a convenient bound for the circumradius. A similar bound in the context of Riemannian manifolds can be found in [40].
Lemma 3.4 (Diameter Estimate).
Let $X$ be a compact $\mathrm{CAT}(\kappa)$ space. Let $A$ be a non-empty, finite subset with $\operatorname{diam}(A)<\Delta(X)$. Then,
  1. (i)
    $A$ has a unique circumcenter $c(A)$ in $X$,
  2. (ii)
    the diameter $\operatorname{diam}\bigl{(}A\bigr{)}\geq\dfrac{4}{3}\operatorname{rad}(A)$.
  3. (iii)
    if $B\subset A$ is a non-empty subset, then $d(c(A),c(B))\leq\dfrac{3}{4}\operatorname{diam}(A)$.
See Appendix for the proof.
3.3. Proofs of Theorems B and C
See B
Proof.
Since $S$ is an $(\xi\beta,\beta)$-approximation of $X$, there exists a correspondence $\mathscr{C}\subset X\times S$ such that for any $(x_{1},s_{1}),(x_{2},s_{2})\in\mathscr{C}$, we have
$\min\{d_{X}(x_{1},x_{2}),d_{S}(s_{1},s_{2})\}\leq\beta\implies|d_{X}(x_{1},x_{2})-d_{S}(s_{1},s_{2})|\leq 2\xi\beta.$ (4)
The correspondence induces (possibly non-continuous and non-unique) maps $\phi:X\to S$ and $\psi:S\to X$ such that $(x,\phi(x))\in\mathscr{C}$ for all $x\in X$ and $(\psi(s),s)\in\mathscr{C}$ for all $s\in S$, respectively. We now justify that the maps on the vertex sets extend to the following simplicial maps of Vietoris–Rips complexes:
$\mathcal{R}_{(1-2\xi)\beta}(X)\xrightarrow{\quad\phi\quad}\mathcal{R}_{\beta}(S)\xrightarrow{\quad\psi\quad}\mathcal{R}_{(1+2\xi)\beta}(X).$ (5)
Simplicial Maps.  To see that $\phi$ is a simplicial map, take an $l$–simplex $\sigma_{l}=[x_{0},x_{1},\ldots,x_{l}]$ in $\mathcal{R}_{(1-2\xi)\beta}(X)$. By definition, we must have $d_{X}(x_{i},x_{j})<(1-2\xi)\beta$ for any $0\leq i,j\leq l$. From (4), we then obtain
$d_{S}(\phi(x_{i}),\phi(x_{j}))\leq d_{X}(x_{i},x_{j})+2\xi\beta<(1-2\xi)\beta+2\xi\beta=\beta.$
Thus the image $\phi(\sigma_{l})=[\phi(x_{0}),\phi(x_{1}),\ldots,\phi(x_{l})]$ is a simplex of $\mathcal{R}_{\beta}(S)$. Similarly, $\psi$ can be shown to be a simplicial map.
The rest of the proof establishes that $\phi$ induces isomorphisms on all homotopy groups. Injectivity.  We argue that the composition $(\psi\circ\phi)$ is contiguous to the natural inclusion: $\mathcal{R}_{(1-2\xi)\beta}(X)\xhookrightarrow{\quad\iota\quad}\mathcal{R}_{(1+2\xi)\beta}(X)$. To prove the claim, take an $l$–simplex $\sigma_{l}=[x_{0},x_{1},\ldots,x_{l}]$ in $\mathcal{R}_{(1-2\xi)\beta}(X)$. Thus, $d_{X}(x_{i},x_{j})<(1-2\xi)\beta$ for all $0\leq i,j\leq l$. We then note from (4) that
$d_{X}((\psi\circ\phi)(x_{i}),x_{j})$ $=d_{X}(\psi(\phi(x_{i})),x_{j})$
$\leq d_{S}(\phi(x_{i}),\phi(x_{j}))+2\xi\beta$
$\leq d_{X}(x_{i},x_{j})+2\xi\beta+2\xi\beta$
$<(1-2\xi)\beta+4\xi\beta=(1+2\xi)\beta.$
This implies that $(\psi\circ\phi)(\sigma_{l})\cup\iota(\sigma_{l})$ is a simplex of $\mathcal{R}_{(1+2\xi)\beta}(X)$. Since $\sigma_{l}$ is an arbitrary simplex, the simplicial maps $(\psi\circ\phi)$ and $\iota$ are contiguous.
Since $(1+2\xi)\beta<\Delta(X)\leq\rho(X)$, Hausmann’s Theorem 3.2 implies that there $\iota$ is a homotopy equivalence. Hence, the induced homomorphism $\iota_{*}$ on the homotopy groups are isomorphisms. On the other hand, we already have $\iota\simeq\psi\circ\phi$. Therefore, the induced homomorphisms $\left(\psi_{*}\circ\phi_{*}\right)$ are also isomorphisms, implying that $\phi_{*}$ is an injective homomorphism on $\pi_{m}\left(\mathcal{R}_{(1-2\xi)\beta}(X)\right)$ for all $m\geq 0$.
Surjectivity.  Next, we argue that $\phi_{*}$ is also a surjective homomorphism on $\pi_{m}\left(\mathcal{R}_{(1-2\xi)\beta}(X)\right)$ for any $m\geq 0$.
As observed in Proposition A.1, both $\mathcal{R}_{(1-2\xi)\beta}(X)$ and $\mathcal{R}_{\beta}(S)$ are path-connected. Thus the result holds for $m=0$.
For $m\geq 1$, let us take an abstract simplicial complex $\mathscr{K}$ such that $\mathscr{K}$ is a triangulation of the $m$-dimensional sphere $\mathbb{S}^{m}$. Note that $\mathscr{K}$ can be assumed to be a pure $m$–complex. In order to show surjectivity of $\phi_{*}$ on $\pi_{m}\left(\mathcal{R}_{(1-2\xi)\beta}(X)\right)$, we start with a simplicial map $g:\mathscr{K}\to\mathcal{R}_{\beta}(S)$ representing any given class in $\pi_{m}\left(\mathcal{R}_{(1-2\xi)\beta}(X)\right)$; such $g$ exists thanks to the Simplicial Approximation Theorem [42, Theorem 16.5]. Next, we argue that there must exist a simplicial map [^9] Here $\operatorname{sd}{\mathscr{K}}=\operatorname{sd}^{1}{\mathscr{K}}$ is a single subdivision of $\mathscr{K}$. $\widetilde{g}:\operatorname{sd}{\mathscr{K}}\to\mathcal{R}_{(1-2\xi)\beta}(X)$ such that the following diagram commutes up to homotopy:
$\mathcal{R}_{(1-2\xi)\beta}(X)$$\mathcal{R}_{\beta}(S)$$\mathscr{K}$$\operatorname{sd}{\mathscr{K}}$$\phi$$g$$\widetilde{g}$$\operatorname{sd}^{-1}$ (6)
where the linear homeomorphism $\operatorname{sd}:\operatorname{sd}{\mathscr{K}}\to\mathscr{K}$ maps each vertex of $\operatorname{sd}{\mathscr{K}}$ to the corresponding point of $\mathscr{K}$. We note that each vertex of $\operatorname{sd}{\mathscr{K}}$ is the barycenter, $\widehat{\sigma}$, of a simplex $\sigma$ of $\mathscr{K}$. To construct the simplicial map $\widetilde{g}:\operatorname{sd}{\mathscr{K}}\to\mathcal{R}_{(1-2\xi)\beta}(X)$, we define it on the vertices of $\operatorname{sd}{\mathscr{K}}$ first, and prove that the vertex map extends to a simplicial map.
Let $\sigma_{l}=[v_{0},v_{1},\ldots,v_{l}]$ be an $l$–simplex of $\mathscr{K}$. Since $g$ is a simplicial map, the image $g(\sigma_{l})=[g(v_{0}),g(v_{1}),\ldots,g(v_{l})]$ is a simplex of $\mathcal{R}_{\beta}(S)$, hence a subset of $S$ with $\operatorname{diam}_{S}(g(\sigma_{l}))<\beta$. For each $0\leq j\leq l$, there exists $x_{j}\in X$ such that $(x_{j},g(v_{j}))\in\mathscr{C}$. We denote by $\sigma_{j}^{\prime}:=[x_{0},x_{1},\ldots,x_{j}]$ for $0\leq j\leq l$. We note from (4), for later, that the diameter of $\sigma_{j}^{\prime}$ is less than $\Delta(X)$:
$\operatorname{diam}_{X}\bigl{(}\sigma_{j}^{\prime}\bigr{)}\leq\operatorname{diam}_{S}\bigl{(}g(\sigma_{j})\bigr{)}+2\xi\beta<\beta+2\xi\beta=(1+2\xi)\beta<\Delta(X).$ (7)
We then define the vertex map
$\widetilde{g}(\widehat{\sigma}_{l})\coloneqq c(\sigma_{l}^{\prime}),$
where $c(\sigma_{l}^{\prime})\in X$ is a circumcenter of $\sigma_{l}^{\prime}$. Due to the diameter bound in (7), Lemma 3.4 implies that a circumcenter of $\sigma_{l}^{\prime}$ exists.
To see that $\widetilde{g}$ extends to a simplicial map, we consider a typical $l$–simplex $\tau_{l}=[\widehat{\sigma}_{0},\ldots,\widehat{\sigma}_{l}]$, of $\operatorname{sd}{\mathscr{K}}$, where $\sigma_{i-1}\prec\sigma_{i}$ for $1\leq i\leq l$ and $\sigma_{i}\in\mathscr{K}$. Now,
$\begin{split}\operatorname{diam}_{X}(\widetilde{g}(\tau_{l}))&=\operatorname{diam}_{X}([c(\sigma_{0}^{\prime}),c(\sigma_{1}^{\prime}),\ldots,c(\sigma_{l}^{\prime}))\\ &=\max_{0\leq i<j\leq l}\{d_{X}(c(\sigma_{i}^{\prime}),c(\sigma_{j}^{\prime}))\}.\end{split}$
Thus,
[…]
Thus $\widetilde{g}(\tau_{l})$ is a simplex of $\mathcal{R}_{(1-2\xi)\beta}(X)$. This implies that $\widetilde{g}$ is a simplicial map.
We lastly invoke Lemma 2.10 to show that the diagram commutes up to homotopy. We need to argue that the simplicial maps $g$ and $(\phi\circ\widetilde{g})$ satisfy the conditions of Lemma 2.10:
  1. (a)
    For any vertex $v\in\mathscr{K}$,
    $(\phi\circ\widetilde{g})(v)=g(v).$
  2. (b)
    For any simplex $\sigma_{m}=[v_{0},v_{1},\ldots,v_{m}]$ of $\mathscr{K}$, we have for $0\leq j\leq m$:
    $d_{S}(g(v_{j}),(\phi\circ\widetilde{g})(\widehat{\sigma}_{m}))$ $=d_{S}(g(v_{j}),\phi(c(\sigma_{m}^{\prime})))$
    $\leq d_{X}\left(x_{j},c(\sigma_{m}^{\prime})\right)+2\xi\beta,\text{ since }(x_{j},g(v_{j}))\in\mathscr{C}$
[…]
    $=\beta-(1-14\xi)\beta/4\leq\beta,\text{ since }\xi\leq\tfrac{1}{14}.$
Thus, $g(\sigma_{m})\cup(\phi\circ\widetilde{g})(\widehat{\sigma}_{m})$ is a simplex of $\mathcal{R}_{\beta}(S)$. Therefore, Lemma 2.10 implies that the diagram commutes. Since $\mathscr{K}=\mathbb{S}^{m}$ and $g$ is arbitrary, we conclude that $\phi$ induces a surjective homomorphism.
Homotopy Equivalence  For any $m\geq 0$, therefore,
$\phi_{*}:\pi_{m}\left(\mathcal{R}_{(1-2\xi)\beta}(X)\right)\to\pi_{m}\left(\mathcal{R}_{\beta}(S)\right).$
is an isomorphism. It follows from Whitehead’s theorem that $\phi$ is a homotopy equivalence. Since $\mathcal{R}_{(1-2\xi)\beta}(X)$ is homotopy equivalent to $X$, we conclude that $\mathcal{R}_{\beta}(S)\simeq X$. ∎
Remark 3.5.
Complexes $\mathcal{R}_{(1-2\xi)\beta}(X)\simeq X$ and $\mathcal{R}_{\beta}(S)$ are a priori infinite complexes, and we know from [41] that Whitehead’s theorem applies for such complexes. On the other hand, $X$ is compact of positive convexity radius $\rho(X)>0$, therefore, the cover by open balls $\{B_{d_{X}}(x,\rho)\}$ is a good cover with a finite subcover. In turn, the Nerve Lemma [10] implies $X$ has a homotopy type of a finite complex, thus Theorems A, B, C, E are concerned with finite complexes up to homotopy.
We immediately obtain our main stability result. See C
Proof.
Treating $X$ or $X^{\prime}$ as the sample metric space $(S,d_{S})$ in Theorem B, we have for instance $\mathcal{R}_{\beta}(X^{\prime})\simeq X$. Since $\beta<\rho(X)$, Haussman’s Theorem 3.2 implies $X^{\prime}\simeq\mathcal{R}_{\beta}(X^{\prime})$. ∎
4. Euclidean Path Metrics and Geodesic Subspaces of $\mathbb{R}^{N}$
This section introduces the notion of the Euclidean path metrics and presents some of their properties relevant to our Euclidean shape reconstruction scheme. In our setting, the underlying hidden shape $X\subset\mathbb{R}^{N}$ is equipped with the Euclidean metric. We denote by $d^{L}_{X}$, the induced length metric on $X$; see Definition 2.2. We consider, for reconstruction, shapes $X\subset\mathbb{R}^{N}$ so that under the induced length metric $(X,d^{L}_{X})$ it is a complete length space—also called a geodesic subspace of $\mathbb{R}^{N}$.
We use the notation $\mathcal{R}^{L}_{\beta}(X)$ to denote the Vietoris–Rips $X$ under the length metric $d^{L}$.
Remark 4.1.
For the Euclidean shape $X$, the metric $(X,d^{L}_{X})$ produces a finer topology on $X$ than $(X,\|\cdot\|)$. The two topologies are not generally the same even when $X$ is compact in the Euclidean topology. One way to force the equivalence is to impose a finite global distortion (Definition 4.2); also see [21, Remark 2.3]. However, the equivalence of the two topologies is not required to build the theory of this paper.
Unless otherwise indicated, we always assume that a geodesic subspace $X\subset\mathbb{R}^{N}$ is compact and path-connected in the $d^{L}_{X}$ topology. As a consequence, so are they in the Euclidean topology.
Definition 4.2 (Global Distortion).
The (global) distortion of embedding to simply distortion (following [29]) is defined as follows
$\delta(X)\coloneqq\displaystyle\sup_{a\neq b\in X}\dfrac{d^{L}_{X}(a,b)}{||a-b||}.$
4.1. The Path Metric and Its Properties
For any $Y\subset\mathbb{R}^{N}$, the notion of an $\varepsilon$–path was introduced in [39], [21] as a family of piecewise paths in $\mathbb{R}^{N}$. The family gives rise to an alternative metric on $Y$—different from the Euclidean submetric.
[…]
Definition 4.3 ($\varepsilon$–Path).
Let $Y\subset\mathbb{R}^{d}$ be non-empty and $\varepsilon>0$ a number. For $a,b\in Y$, an $\varepsilon$–path of $Y$ from $a$ to $b$ is a finite sequence $P=\{y_{i}\}_{i=0}^{k+1}\subseteq Y$ such that $y_{0}=a$, $y_{k+1}=b$, and $\left\lVert y_{i}-y_{i+1}\right\rVert<\varepsilon$ for all $i=0,1,\ldots,k$.
We now define the length of the path by
$L(P)\coloneqq\sum_{i=0}^{k}\left\lVert y_{i}-y_{i+1}\right\rVert.$
We denote the set of all $\varepsilon$–paths of $Y$ by $\mathscr{P}_{Y}^{\varepsilon}$. We finally introduce the $\varepsilon$–path metric on $Y$.
Definition 4.4 ($d^{\varepsilon}_{Y}$–Metric).
Let $Y\subset\mathbb{R}^{d}$ be non-empty and $\varepsilon>0$ a number. The $\varepsilon$–path metric on $Y$, denoted $d^{\varepsilon}_{Y}$, between any $a,b\in Y$ is defined by
$d^{\varepsilon}_{Y}(a,b)\coloneqq\inf\left\{L(P)\mid P\in\mathscr{P}^{\varepsilon}_{Y}\text{ is an $\varepsilon$--path of $Y$ between $a$ and $b$}\right\}.$
We remark that $d^{\varepsilon}_{Y}(a,b)$ is not finite, in general, for all $\varepsilon>0$. For it to be finite, the Euclidean triangle inequality suggests that the Euclidean thickening $Y^{\frac{\varepsilon}{2}}$ of $Y$ must be path-connected. However, when $d^{\varepsilon}_{Y}$ is finite, $(Y,d^{\varepsilon}_{Y})$ is a metric space as noted in the following proposition.
Proposition 4.5 (Path-metric Space).
Let $Y\subset\mathbb{R}^{N}$ and $\varepsilon>0$ be a number such that the Euclidean thickening $Y^{\frac{\varepsilon}{2}}$ is path-connected. Then, $(Y,d^{\varepsilon}_{Y})$ is a metric space.
Proof.
Let $a,b\in Y$. We show that there must exist an $\varepsilon$–path of $Y$ joining them. Since $Y^{\frac{\varepsilon}{2}}=\cup_{y\in Y}\mathbb{B}(y,\frac{\varepsilon}{2})$ is path-connected, there exists a sequence of points $\{y_{i}\}_{i=0}^{k+1}\subset Y$ such that $y_{0}=a$, $y_{k+1}=b$, and $\mathbb{B}(y_{i},\frac{\varepsilon}{2})\cap\mathbb{B}(y_{i+1},\frac{\varepsilon}{2})$ is non-empty for each $i\in\{1,\ldots,k\}$. As a consequence, $\left\lVert y_{i}-y_{i+1}\right\rVert<\varepsilon$ for each $i\in\{1,\ldots,k\}$. Thus, $\{y_{i}\}_{i=0}^{k}$ is indeed an $\varepsilon$–path between $a$ and $b$. ∎
In the metric space $(Y,d^{\varepsilon}_{Y})$, we denote the diameter of a subset $A\subset Y$ by $\operatorname{diam}_{\varepsilon}(A)$. For any scale $\beta>0$, the Vietoris–Rips complex of $(Y,d^{\varepsilon}_{Y})$ is denoted by $\mathcal{R}^{\varepsilon}_{\beta}(Y)$.
We now present some interesting properties of path-metrics.
Proposition 4.6 (Non-decreasing).
Let $Y\subset\mathbb{R}^{N}$ be a subset and $a,b\in Y$. Then, $d_{Y}^{\varepsilon}(a,b)$ is a non-increasing function of $\varepsilon$.
Proof.
Let $\varepsilon_{1}<\varepsilon_{2}$ and $0<\eta<\varepsilon_{2}-\varepsilon_{1}$ be arbitrary. By the definition of $\varepsilon$–paths, there exists an $\varepsilon_{1}$–path $P$ joining $a$ and $b$ such that $L(P)\leq d_{Y}^{\varepsilon_{1}}(a,b)+\eta$. Since $\varepsilon_{1}<\varepsilon_{2}$, note that $P$ is also an $\varepsilon_{2}$–path joining $a,b$. Thus,
$d_{Y}^{\varepsilon_{2}}(a,b)\leq L(P)\leq d_{Y}^{\varepsilon_{1}}(a,b)+\eta.$
Since $\eta$ is arbitrary, we conclude that $d_{Y}^{\varepsilon_{2}}(a,b)\leq d_{Y}^{\varepsilon_{1}}(a,b)$. ∎
The following Proposition from [39, Prop. 4.6] is of central importance.
Proposition 4.7 (Comparison of Path Metrics).
Let $X\subset\mathbb{R}^{N}$ be a geodesic subspace and $S\subset\mathbb{R}^{N}$ such that $d_{H}(S,X)<\frac{1}{2}\xi\varepsilon$ for some $\xi\in(0,1)$ and $\varepsilon>0$. For any $x_{1},x_{2}\in X$ and corresponding $s_{1},s_{2}\in S$ with $||x_{1}-s_{1}||,||x_{2}-s_{2}||<\frac{1}{2}\xi\varepsilon$, we have
$\left\lVert s_{1}-s_{2}\right\rVert\leq d^{\varepsilon}_{S}(s_{1},s_{2})\leq\dfrac{d^{L}_{X}(x_{1},x_{2})+\xi\varepsilon}{1-\xi}.$ (8)
Note the first inequality comes from the triangle inequality, whereas the second inequality directly follows from [39, Prop. 4.6].
4.2. Stability of Path Metrics
The following result demonstrates the stability of the path metric under Hausdorff perturbation. Although we do not directly apply it in later arguments, we record it here as of independent interest.
Theorem 4.8 (Stability of Path Metrics).
Let $X,S\subset\mathbb{R}^{N}$ be any subsets with $d_{H}(X,S)<\frac{1}{2}\xi\varepsilon$ for some $\xi\in(0,1)$ and $\varepsilon>0$. Let $p,q\in S$ and $p^{\prime},q^{\prime}\in X$ such that $\left\lVert p-p^{\prime}\right\rVert<\frac{1}{2}\xi\varepsilon$ and $\left\lVert q-q^{\prime}\right\rVert<\frac{1}{2}\xi\varepsilon$. Then, $d^{(2+\xi)\varepsilon}_{X}(p^{\prime},q^{\prime})\leq(1+\xi)d^{\varepsilon}_{S}(p,q)$.
Proof.
The result trivially follows if $d^{\varepsilon}_{S}(p,q)<\varepsilon$. Thus, we assume that $d^{\varepsilon}_{S}(p,q)\geq\varepsilon$.
Let $P=\{x_{i}\}_{i=0}^{l+1}$ be an $\varepsilon$–path in $S$ joining $p,q$, i.e. $x_{0}=p$ and $x_{l+1}=q$. Since $d^{\varepsilon}_{S}(p,q)\geq\varepsilon$, there must exist a subsequence $\{x_{i_{j}}\}_{j=0}^{k+1}$ of $P$ such that $i_{0}=0$, $i_{k+1}=l+1$, and
$\varepsilon\leq\sum_{i=i_{j}}^{i_{j+1}}\left\lVert x_{i}-x_{i+1}\right\rVert<2\varepsilon\text{ for any }0\leq j\leq k.$ (9)
As a consequence, the total length
$L(P)=\sum_{i=0}^{l+1}\left\lVert x_{i}-x_{i+1}\right\rVert=\sum_{j=0}^{k}\sum_{i=i_{j}}^{i_{j+1}}\left\lVert x_{i}-x_{i+1}\right\rVert\geq\sum_{j=0}^{k}\varepsilon=(k+1)\varepsilon.$
Now, for each $j$, there is a corresponding point $x_{i_{j}}^{\prime}\in X$ such that $\left\lVert x_{i_{j}}-x^{\prime}_{i_{j}}\right\rVert<\frac{1}{2}\xi\varepsilon$. We choose $x^{\prime}_{i_{0}}=p^{\prime}$ and $x^{\prime}_{i_{k+1}}=q^{\prime}$. For any $0\leq j\leq k$, we observe first from the triangle inequality and afterward from (9) that
$\left\lVert x^{\prime}_{i_{j}}-x^{\prime}_{i_{j+1}}\right\rVert\leq\left\lVert x_{i_{j}}-x_{i_{j+1}}\right\rVert+\xi\varepsilon\leq\sum_{i=i_{j}}^{i_{j+1}}\left\lVert x_{i}-x_{i+1}\right\rVert+\xi\varepsilon<2\varepsilon+\xi\varepsilon=(2+\xi)\varepsilon.$
Thus, the sequence of points $P^{\prime}=\{x^{\prime}_{i_{j}}\}_{j=0}^{k+1}$ forms a $(2+\xi)\varepsilon$–path in $X$ joining $p^{\prime}$ and $q^{\prime}$. Moreover, its length
$L(P^{\prime})$ $=\sum_{j=0}^{k}\left\lVert x^{\prime}_{i_{j}}-x^{\prime}_{i_{j+1}}\right\rVert\leq\sum_{j=0}^{k}\left(\left\lVert x_{i_{j}}-x_{i_{j+1}}\right\rVert+\xi\varepsilon\right)\leq\sum_{j=0}^{k}\left(\sum_{i=i_{j}}^{i_{j+1}}\left\lVert x_{i}-x_{i+1}\right\rVert+\xi\varepsilon\right)$
$=L(P)+\sum_{j=0}^{k}\xi\varepsilon=L(P)+(k+1)\xi\varepsilon\leq L(P)+\xi L(P)=(1+\xi)L(P).$
By the definition of path metric, we can write $d_{X}^{(2+\xi)\varepsilon}(p^{\prime},q^{\prime})\leq L(P^{\prime})\leq(1+\xi)L(P)$. This is true for any arbitrary $\varepsilon$–path $P$ joining $p$ and $q$. Therefore, we conclude that
$d^{(2+\xi)\varepsilon}_{X}(p^{\prime},q^{\prime})\leq(1+\xi)d^{\varepsilon}_{S}(p,q).\qed$
4.3. Convergence of Path Metrics
Intuitively, as $\varepsilon$ go to 00, the length of the $\varepsilon$–path between two points $x_{1},x_{2}\in X$ converges to the length of the geodesic path joining them. We make the intuition more precise in the following result, which is reminiscent of Gromov–Hausdorff convergence Theorem in [12, p. 265, Theorem 7.5.1] in the specific setting of thickenings $X^{\varepsilon}$ and Hausdorff closeness.
Theorem 4.9 (Convergence of Path Metric).
Let $X\subset\mathbb{R}^{N}$ be a geodesic subspace. Then, $d^{\varepsilon}_{X}$ converges to $d^{L}_{X}$ uniformly on $X\times X$ as $\varepsilon\to 0$.
Recall $X^{\varepsilon}$ for any $\varepsilon>0$, is the $\varepsilon$–thickening of $X$ in $\mathbb{R}^{N}$. We denote by $d^{L}_{X^{\varepsilon}}$ the induced length metric on $X^{\varepsilon}$. Moreover, for any $x_{1},x_{2}\in X$, we must have that $d^{L}_{X^{\varepsilon}}(x_{1},x_{2})\leq d^{L}_{X}(x_{1},x_{2})$.
Theorem 4.9 follows almost immediately from the following technical lemma concerning the convergence of induced length metrics.
Lemma 4.10 (Convergence of Induced Length Metric).
Let $X\subset\mathbb{R}^{N}$ be a geodesic subspace and $x_{1},x_{2}\in X$. Then, $d^{L}_{X^{\varepsilon}}(x_{1},x_{2})\leq d^{L}_{X}(x_{1},x_{2})$, and
$d^{L}_{X^{\varepsilon}}(x_{1},x_{2})\longrightarrow d^{L}_{X}(x_{1},x_{2})\text{ as }\varepsilon\to 0.$
Proof of Lemma 4.10.
For each small $\varepsilon>0$, let $\gamma_{\varepsilon}$ be a path in $X^{\varepsilon}$ connecting $x_{1}$ and $x_{2}$ such that $L(\gamma_{\varepsilon})\leq d^{L}_{X^{\varepsilon}}(x_{1},x_{2})+\varepsilon$. Consequently, the lengths of the paths are bounded above; let $L$ be an upper bound, i.e., $L\geq L(\gamma_{\varepsilon})$ for all $\varepsilon$.
Since $X$ is a compact length space, the sequence $\{\gamma_{\varepsilon}\}$ is uniformly bounded. Using the diagonal argument (as used in the proof of Arzela-Ascolli’s them [47]), we obtain a subsequence $\{\gamma_{\varepsilon_{i}}\}$ pointwise convergent to a function $\gamma:I\longrightarrow X$. We now argue that the convergence is uniform and that $L(\gamma_{\varepsilon_{i}})\longrightarrow L(\gamma)$.
We (re)-parameterize each $\gamma_{\varepsilon_{i}}(t)$ with the arc-length parameter $t_{i}=\phi_{i}(s)$ as follows:
$\begin{split}\gamma_{i}&\coloneqq\gamma_{\varepsilon_{i}}:[0,L]\longrightarrow X^{2\varepsilon},\ \gamma_{i}(0)=x_{1},\\ \gamma_{i}(s)&=\gamma_{\varepsilon_{i}}(\phi_{i}(s)),\quad s\in[0,L(\gamma_{i})],\\ \gamma_{i}(s)&=x_{2},\quad s\in[L(\gamma_{i}),L].\end{split}$
Let $\overline{\varepsilon}>0$ be any small number. Then, we can choose $n\in\mathbb{N}$ such that $\frac{L}{n}<\frac{1}{3}\overline{\varepsilon}$. We then subdivide $[0,L]$ in equal intervals:
$0=s_{0}<s_{1}<\ldots<s_{n}=L\text{ and }s_{j}=s_{j-1}+\frac{L}{n}=j\frac{L}{n}\text{ for }1\leq j\leq n.$
Thanks to the pointwise convergence for any given $\overline{\varepsilon}>0$, there exists an $N(\overline{\varepsilon})\in\mathbb{N}$, such that for every $k,m>N(\overline{\varepsilon})$:
$\|\gamma_{k}(s_{j})-\gamma_{m}(s_{j})\|<\tfrac{1}{3}\overline{\varepsilon},\qquad\text{for every}\quad j=0,\ldots,n.$ (10)
(i.e. Cauchy’s condition holds at all points in $\{s_{j}\}_{j=0}^{n}$). We note now that the condition extends to any $s\in[s_{j},s_{j+1}]$:
$\|\gamma_{k}(s)-\gamma_{m}(s)\|$ $\leq\|\gamma_{k}(s)-\gamma_{k}(s_{j})\|+\|\gamma_{k}(s_{j})-\gamma_{m}(s_{j})\|+\|\gamma_{m}(s_{j})-\gamma_{m}(s)\|$
$<L\left(\gamma_{k}\big{|}_{[s_{j},s_{j+1}]}\right)+\tfrac{\overline{\varepsilon}}{3}+L\left(\gamma_{m}\big{|}_{[s_{j},s_{j+1}]}\right),\text{ since Cauchy at }s_{j}$
[…]
$<\tfrac{\overline{\varepsilon}}{3}+\tfrac{\overline{\varepsilon}}{3}+\tfrac{\overline{\varepsilon}}{3}=\overline{\varepsilon}.$
This proves that the sequence $\{\gamma_{k}\}$ is uniformly Cauchy, and thus uniformly convergent to $\gamma$. Consequently, $\gamma$ is a continuous path connecting $x_{1}$ and $x_{2}$.
Applying Lemma 2.1 on $(X^{\varepsilon},\|\cdot\|)$, we get from the semi-continuity of the induced length
$\lim\inf L(\gamma_{\varepsilon_{i}})\geq L(\gamma).$
As a result, the length $L(\gamma)\leq\liminf L(\gamma_{\varepsilon_{i}})\leq\liminf[d^{L}_{X^{\varepsilon}}(x_{1},x_{2})+\varepsilon]$. On the other hand, $d^{L}_{X^{\varepsilon}}(x_{1},x_{2})\leq d^{L}_{X}(x_{1},x_{2})\leq L(\gamma)$. Thus, we have
$\limsup\limits_{\varepsilon\to 0}d^{L}_{X^{\varepsilon}}(x_{1},x_{2})\leq L(\gamma)\leq\liminf\limits_{\varepsilon\to 0}[d^{L}_{X^{\varepsilon}}(x_{1},x_{2})+\varepsilon].$
Therefore, the limit exists: $L(\gamma)=\lim d^{L}_{X^{\varepsilon}}(x_{1},x_{2})$. ∎
Now, we prove the convergence result.
Proof of Theorem 4.9.
For any $x_{1},x_{2}\in X$, the pointwise convergence follows from Lemma 4.10 by sending $\varepsilon\to 0$ in
$d^{L}_{X^{2\varepsilon}}(x_{1},x_{2})\leq d^{\varepsilon}_{X}(x_{1},x_{2})\leq d^{L}_{X}(x_{1},x_{2}).$
Here, the first inequality comes from the fact that an $\varepsilon$-path on $X$ is a continuous path joining $x_{1}$ and $x_{2}$ in $X^{2\varepsilon}$, and the second inequality follows from the fact that we may always choose an $\varepsilon$-path along the shortest path connecting $x_{1}$ and $x_{2}$.
We note that $d^{\varepsilon}_{X}$ is a monotone non-decreasing function of $\varepsilon$ (as $\varepsilon$ decreases). Moreover, both $d^{L}_{X}$ and $d^{\varepsilon}_{X}$ are continuous w.r.t. the $d^{L}$ topology on $X\times X$. Now, the uniform convergence follows from Dini’s theorem (see [8] for example) and the assumption that $X\times X$ is compact w.r.t. the $d^{L}_{X}$ topology. ∎
As an immediate consequence, we note the following Gromov–Hausdorff convergence.
Proposition 4.11.
For a geodesic subspace $X\subset\mathbb{R}^{N}$, the path metric $(X,d^{\varepsilon}_{X})$ converges to the induced length metric $(X,d^{L}_{X})$ in the Gromov–Hausdorff distance as $\varepsilon\to 0$.
Proof.
Fix $\overline{\varepsilon}>0$. From the uniform convergence in Theorem 4.9, we can choose an $\varepsilon>0$ such that
$\sup_{x_{1},x_{2}\in X}|d^{\varepsilon}_{X}(x_{1},x_{2})-d^{L}_{X}(x_{1},x_{2})|\leq\overline{\varepsilon}$
Now, choosing the diagonal correspondence between $X$ and itself, note from its (Gromov–Hausdorff) distortion that
$d_{GH}((X,d^{\varepsilon}_{X}),(X,d^{L}_{X}))\leq\frac{1}{2}\sup_{x_{1},x_{2}\in X}|d^{\varepsilon}_{X}(x_{1},x_{2})-d^{L}_{X}(x_{1},x_{2})|\leq\overline{\varepsilon}.$
This concludes the result. ∎
5. Reconstruction of Euclidean Subspaces under Hausdorff Distance
We define our most important sampling parameter: restricted distortion.
Let $X\subset\mathbb{R}^{N}$ be a geodesic subspace. The restricted distortion of $X$, denoted $\delta^{\varepsilon}_{R}(X)$, is parametrized by $\varepsilon>0$ and $R>0$. Roughly speaking, it is supremum the ratio of the induced length of metrics on $X$ and (Euclidean) $\varepsilon$-thickening of $X$, between points of $X$ that are at least $R$ distance away. More formally, we present the following definition.
Definition 5.1 (Restricted Distortion).
For $\varepsilon>0$ and $R>0$, the restricted distortion or $(\varepsilon,R)$-distortion of a geodesic subspace $X\subset\mathbb{R}^{N}$ is defined as
$\delta^{\varepsilon}_{R}(X)\coloneqq\sup_{d^{L}_{X}(x_{1},x_{2})\geq R}\frac{d_{X}^{L}(x_{1},x_{2})}{d^{L}_{X^{\varepsilon}}(x_{1},x_{2})}.$ (11)
It immediately follows from the definition that $\delta^{\varepsilon}_{R}(X)$ is a non-decreasing function of $\varepsilon$ and that $\delta^{\varepsilon}_{R}(X)$ is a non-increasing function of $R$. Moreover, $\delta^{\varepsilon}_{R}(X)\leq\delta(X)$, the global distortion of embedding of $X$. Indeed, we have
$\lim_{\begin{subarray}{c}\varepsilon\to\infty\\ R\to 0\end{subarray}}\delta^{\varepsilon}_{R}(X)=\delta(X).$
Remark 5.2.
A slightly modified version of Definition 5.1 given as
$\widehat{\delta}^{\varepsilon}_{R}(X)\coloneqq\sup_{d^{L}_{X}(x_{1},x_{2})\geq R}\frac{d_{X}^{L}(x_{1},x_{2})}{d^{\varepsilon}_{X}(x_{1},x_{2})}$
can be also applied in Theorem A and E. This version of restricted distortion appears more computationally amenable. The metric properties, estimates for relevant classes of spaces (metric graphs and manifolds), and applications of $\widehat{\delta}^{\varepsilon}_{R}(X)$ are further investigated in [34] and [27].
Now, we present the most important convergence property of restricted distortion.
Proposition 5.3 (Convergence of Restricted Distoriton).
Let $X$ be a geodesic subspace of $\mathbb{R}^{N}$. For any $R>0$, we have $\delta^{\varepsilon}_{R}(X)\to 1$ as $\varepsilon\to 0$.
Proof.
Let $x_{1},x_{2}\in X$ be such that $x_{1}\neq x_{2}$. We have already established in Theorem 4.9 that the ratio
$\delta^{\varepsilon}(x_{1},x_{2})\coloneqq\frac{d^{L}(x_{1},x_{2})}{d^{L}_{X^{\varepsilon}}(x_{1},x_{2})}\to 1\text{ as }\varepsilon\to 0.$
We argue that $\delta^{\varepsilon}(x_{1},x_{2})$ is also continuous at $x_{1},x_{2}\in X$ w.r.t. $(X,d^{L})$. The numerator, length metric $d^{L}_{X}:X\times X\to\mathbb{R}$, is naturally continuous with respect to the $(X,d^{L})$. Similarly, the denominator $d^{L}_{X^{\varepsilon}}\colon X\times X\to\mathbb{R}$ is continuous w.r.t $(X,d^{L}_{X^{\varepsilon}})$, the subspace topology of the induced length metric on $X^{\varepsilon}$. Since $d^{L}_{X^{\varepsilon}}(x_{1},x_{2})\leq d^{L}_{X}(x_{1},x_{2})$, the latter topology $X$ in finer. Consequently, the denominator $d^{L}_{X^{\varepsilon}}\colon X\times X\to\mathbb{R}$ remains continuous when assumed the latter topology $(X,d^{L}_{X})$ on the domain. Next, note that $d^{L}_{X^{\varepsilon}}$ is bounded away from zero when restricted to
$X^{2}_{R}=X\times X-\Delta^{\varepsilon}_{R}=\left\{(x_{1},x_{2})\in X\times X\mid\ d^{L}_{X}(x_{1},x_{2})\geq R\right\},$
indeed, from compactness of $X^{2}_{R}$, $\min_{X^{2}_{R}}d^{L}_{X^{\varepsilon}}$ is achieved at some $(a_{0},b_{0})\in X^{2}_{R}$, if $d^{L}_{X^{\varepsilon}}(a_{0},b_{0})=0$, then $a_{0}=b_{0}$ which yields a contradiction. As a result, the ratio $\delta^{\varepsilon}$ is continuous on $X^{2}_{R}$. Since we have pointwise convergence of $\{\delta^{\varepsilon}:X^{2}_{R}\to\mathbb{R}\}$ to $1$, and it is a non-descreasing function of $\varepsilon$, Dini’s theorem (see [8] for example) implies uniform convergence. ∎
We prove the following technical result.
Proposition 5.4 (Restricted Distortion and Path Metrics).
Let $X$ be a geodesic subspace of $\mathbb{R}^{N}$ and $S\subset\mathbb{R}^{N}$ compact with $d_{H}(X,S)<\alpha$. Let $\varepsilon>2\alpha$ and $R$ be positive numbers. For any $x_{1},x_{2}\in X$ such that $d^{L}_{X}(x_{1},x_{2})\geq R$ and $s_{1},s_{2}\in S$ such that $\|x_{i}-s_{i}\|<\alpha$ for $i=1,2$, we must have
$d^{L}_{X}(x_{1},x_{2})\leq\delta^{\varepsilon}_{R}(X)(d^{\varepsilon}_{S}(s_{1},s_{2})+2\alpha).$
Proof.
We consider the path $\gamma:[a,b]\to\mathbb{R}^{N}$ joining $x_{1}$ and $x_{2}$ by concatenating three pieces: the segment $[x_{1},s_{1}]$ (of length $\leq\alpha$) with an $\varepsilon$–path connecting $s_{1}$ and $s_{2}$ (of length $d^{\varepsilon}_{S}(s_{1},s_{2})$), and the segment $[s_{2},x_{2})]$ (of length $\leq\alpha$).
Since $\alpha<\varepsilon/2$, it’s not difficult to see that the image of $\gamma$ lies entirely in $X^{\varepsilon}$. This implies that $d^{L}_{X^{\varepsilon}}(x_{1},x_{2})\leq L(\gamma)\leq(d^{\varepsilon}_{S}(s_{1},s_{2})+2\alpha)$.
From the definition of restricted distortion, we now get
$d^{L}_{X}(x_{1},x_{2})\leq\delta^{\varepsilon}_{R}(X)d^{L}_{X^{\varepsilon}}(x_{1},x_{2})\leq\delta^{\varepsilon}_{R}(X)(d^{\varepsilon}_{S}(s_{1},s_{2})+2\alpha).\qed$
5.1. Proofs of Theorems E and A
See E
Proof.
Let $S\subset\mathbb{R}^{N}$ be compact with $d_{H}(X,S)<\frac{1}{2}\xi\varepsilon$. We define the following relation $\mathscr{C}\subset X\times S$:
$\mathscr{C}=\left\{(x,s)\in X\times S\colon\|x-s\|<\tfrac{1}{2}\xi\varepsilon\right\}.$
Since $d_{H}(X,S)<\frac{1}{2}\xi\varepsilon$, the relation $\mathscr{C}$ is, in fact, a correspondence.
Take arbitrary $(x_{1},s_{1}),(x_{2},s_{2})\in\mathscr{C}$ such that $\min\{d_{X}(x_{1},x_{2}),d_{S}(s_{1},s_{2})\}\leq\beta$. Without any loss of generality, we further assume that $d_{S}(s_{1},s_{2})\leq\beta$. We now argue that:
$\left\lvert d^{L}_{X}(x_{1},x_{2})-d^{\varepsilon}_{S}(s_{1},s_{2})\right\rvert\leq 2\xi\beta.$
We consider the following two cases:
  1. (1)
    If $d^{L}_{X}(x_{1},x_{2})\leq d^{\varepsilon}_{S}(s_{1},s_{2})$, then from the right inequality of Proposition 4.7 we get
    $\left\lvert d^{L}_{X}(x_{1},x_{2})-d^{\varepsilon}_{S}(s_{1},s_{2})\right\rvert$ $=d^{\varepsilon}_{S}(s_{1},s_{2})-d^{L}_{X}(x_{1},x_{2})\leq d^{\varepsilon}_{S}(s_{1},s_{2})-[(1-\xi)d^{\varepsilon}_{S}(s_{1},s_{2})-\xi\varepsilon]$
    $=\xi d^{\varepsilon}_{S}(s_{1},s_{2})+\xi\varepsilon\leq 2\xi\beta.$
    The last inequality is due to the fact that $\varepsilon\leq\beta$ and $d^{\varepsilon}_{S}(s_{1},s_{2})\leq\beta$.
  2. (2)
    If $d^{L}_{X}(x_{1},x_{2})\geq d^{\varepsilon}_{S}(s_{1},s_{2})$, then $\left\lvert d^{L}_{X}(x_{1},x_{2})-d^{\varepsilon}_{S}(s_{1},s_{2})\right\rvert=d^{L}_{X}(x_{1},x_{2})-d^{\varepsilon}_{S}(s_{1},s_{2})$. We consider the following two sub-cases:
    If $d^{L}_{X}(x_{1},x_{2})<2\xi\beta$, then
    $\left\lvert d^{L}_{X}(x_{1},x_{2})-d^{\varepsilon}_{S}(s_{1},s_{2})\right\rvert=d^{L}_{X}(x_{1},x_{2})-d^{\varepsilon}_{S}(s_{1},s_{2})\leq d^{L}_{X}(x_{1},x_{2})<2\xi\beta.$
    If $d^{L}_{X}(x_{1},x_{2})\geq 2\xi\beta$, then Proposition 5.4 implies that
    $d^{L}_{X}(x_{1},x_{2})\leq\delta^{\varepsilon}_{2\xi\beta}(X)(d^{\varepsilon}_{S}(s_{1},s_{2})+\xi\varepsilon).$
    As a result,
    $\left\lvert d^{L}_{X}(x_{1},x_{2})-d^{\varepsilon}_{X}(s_{1},s_{2})\right\rvert$ $=d^{L}_{X}(x_{1},x_{2})-d^{\varepsilon}_{S}(s_{1},s_{2})$
    $\leq\delta^{\varepsilon}_{2\xi\beta}(X)(d^{\varepsilon}_{S}(s_{1},s_{2})+\xi\varepsilon)-d^{\varepsilon}_{S}(s_{1},s_{2})$
    $\leq\tfrac{1+2\xi}{1+\xi}(d^{\varepsilon}_{S}(s_{1},s_{2})+\xi\varepsilon)-d^{\varepsilon}_{S}(s_{1},s_{2})$
    $=\tfrac{\xi}{1+\xi}d^{\varepsilon}_{S}(s_{1},s_{2})+\tfrac{\xi(1+2\xi)}{1+\xi}\varepsilon$
    $\leq\tfrac{\xi}{1+\xi}\beta+\tfrac{\xi(1+2\xi)}{1+\xi}\beta\text{since }d^{\varepsilon}_{S}(s_{1},s_{2})\leq\beta\text{ and }\varepsilon\leq\beta$
    $=2\xi\beta.$
Combining all the cases, we conclude that
$\left\lvert d^{L}_{X}(x_{1},x_{2})-d^{\varepsilon}_{S}(s_{1},s_{2})\right\rvert\leq 2\xi\beta.$
Hence, $(S,d^{\varepsilon}_{S})$ is an $(\xi\beta,\beta)$-approximation of $(X,d^{L})$. ∎
Combining Theorems B and E yields our main result:
See A
6. More on Distortion and $\mu$–reach
In this section, we compare the well-known sampling parameter $r_{\mu}(X)$, the $\mu$–reach of a compact subset $X\subset\mathbb{R}^{N}$ (see Chazal, Cohen-Steiner, and Lieutier [14]) and the restricted distortion parameter $\delta^{\varepsilon}_{R}(X)$ used in our main results above. We intend to show an example that favors our restricted distortion parameter over the $\mu$–reach, in case $X$ is a $\mathrm{CAT}(\kappa)$ subspace of $\mathbb{R}^{N}$.
Let us denote by $R_{X}:\mathbb{R}^{N}\rightarrow[0,\infty)$ the Euclidean distance function defined as $R_{X}(a)=\min_{x\in X}\|a-x\|$. Recall from [14] that the generalized gradient function of $R_{X}$ is defined by
$\nabla_{X}(x)=\dfrac{x-\Theta_{X}(x)}{R_{X}(x)},$ (12)
where $\Theta_{X}(x)$ is the point closest to $x$ in the convex hull of the set:
$\Gamma_{X}(x)=\{y\in X\mid||x-y||=R_{X}(x)\}.$ (13)
It follows that $x_{0}\in\mathbb{R}^{N}$ is a critical point of $R_{X}$ (i.e. $\nabla_{X}(x_{0})=0$) if and only if $x_{0}$ is in $\Gamma_{K}(x_{0})$; in the case $X$ is a submanifold, $R_{X}$ is differentiable and $\nabla_{X}$ is the same as the classical gradient. Next, the critical function $\chi_{X}:(0,\infty)\rightarrow\mathbb{R}^{+}$ of $X$ is defined as
$\chi_{X}(d)=\displaystyle\inf_{y\in R^{-1}_{X}(d)}||\nabla_{X}(y)||$ (14)
i.e. for each fixed value of the distance to $X$, $\chi_{X}(d)$ is the infimum of $\|\nabla_{X}(\,\cdot\,)\|$ along the $d$–level set of $R_{X}$. Finally, the $\mu$-reach of a compact subset $X\subset\mathbb{R}^{N}$ is defined to be
$r_{\mu}(X)=\displaystyle\inf\{d\mid\chi_{X}(d)<\mu\}.$ (15)
Figure 1. (left) Star–shaped subspace $\mathcal{S}$ (Ninja’s Star) (right) $r_{\mu}(\mathcal{S})=0$ at the cusp.
Consider the cusped star–shaped subspace $\mathcal{S}$ of $\mathbb{R}^{2}$ as shown in Figure 1 (left) which is a union of four-quarter circular arcs. Let $z$ be a point near the cusp, which sits on the bisector line of the cusp, within distance $d$ to $\mathcal{S}$. From (12) and Figure 1 (right), we obtain $\Gamma_{\mathcal{S}}(z)=\{a_{0},a_{1}\}$ and $\Theta(z)$ is a midpoint of the segment $[a_{0},a_{1}]$, thus
$\nabla_{\mathcal{S}}(z)=\frac{z-\Theta(z)}{d},\quad\text{and}\quad\|\nabla_{\mathcal{S}}(z)\|=|\sin(\varphi)|.$
Clearly, as $z$ approaches the cusp point $z_{0}$, we have $d\longrightarrow 0$ and $\varphi\longrightarrow 0$, thus $\|\nabla_{\mathcal{S}}(z)\|\longrightarrow 0$. Applying definitions (14) and (15) we conclude that for any value of $\mu$, the $\mu$–reach $r_{\mu}(\mathcal{S})$ vanishes. On the other hand, we clearly have $0<\delta^{\varepsilon}_{R}(\mathcal{S})<\infty$ for any value of $R>0$ and $\varepsilon>0$, also the intrinsic convexity radius of $\mathcal{S}$ is positive and the curvature is negative. Therefore, Theorem A is applicable in the case of $X=\mathcal{S}$ or other similar “cuspy” shapes. In the context of the previous works [21], note that $\delta(X)$ is infinite at the cups of $\mathcal{S}$ (or other types of cusps as well), see [50] for alternative approaches to this issue and the concept of $\alpha$–distortion.
Appendix A Additional Proofs
Proof of Lemma 3.4.
First note that the closed ball $\overline{\mathbb{B}(a,\operatorname{diam}(A))}$ for any $a\in A$ contains $A$ entirely, implying that $\operatorname{rad}(A)\leq\operatorname{diam}(A)$. Since $\operatorname{diam}(A)<\Delta(X)\leq\frac{\pi}{4\sqrt{\kappa}}$, we therefore have $\operatorname{rad}(A)\in\left[0,\frac{\pi}{4\sqrt{\kappa}}\right]$. Thus, Part (i) follows immediately from Jung’s Theorem 3.3.
[…]
Case 1 ($\kappa\leq 0$)
From (3), we get
$\operatorname{diam}(A)\geq 2\operatorname{rad}(A)\sqrt{\frac{n+1}{2n}}\geq 2\operatorname{rad}(A)\frac{1}{\sqrt{2}}\geq\frac{4}{3}\operatorname{rad}(A).$
Case 2 ($\kappa>0$)
Let us now define the function $J(r)\coloneqq\frac{2}{\sqrt{\kappa}}\sin^{-1}\left(\sqrt{\frac{n+1}{2n}}\sin{\sqrt{\kappa}r}\right)$, and show that $J(r)/r$ is a strictly decreasing function for $r\in\left(0,\frac{\pi}{4\sqrt{\kappa}}\right)$. We take its derivative:
$\frac{d}{dr}(J(r)/r)=\frac{rJ^{\prime}(r)-J(r)}{r^{2}}=\frac{f(r)}{r^{2}},$ (16)
where $f(r)\coloneqq rJ^{\prime}(r)-J(r)$. We observe that $f(r)$ is strictly decreasing, since
$f^{\prime}(r)=rJ^{\prime\prime}(r)=r\frac{2\sqrt{\kappa}\sqrt{\frac{n+1}{2n}}\sin{\sqrt{\kappa}r}}{\left(1-\frac{n+1}{2n}\sin^{2}{\sqrt{\kappa}r}\right)^{3/2}}\left(\frac{1-n}{2n}\right)<0$
for $r\in\left(0,\frac{\pi}{4\sqrt{\kappa}}\right)$ and $n\geq 2$. Then it follows from $f(0)=0$ that $f(r)<0$ on $r\in\left(0,\frac{\pi}{4\sqrt{\kappa}}\right)$.
Thus, (16) implies that the function $J(r)/r$ is a strictly decreasing function for $r\in\left(0,\frac{\pi}{4\sqrt{\kappa}}\right)$. Its minimum, as a result, is attained at $r=\frac{\pi}{4\sqrt{\kappa}}$. Thus, for $r\in\left(0,\frac{\pi}{4\sqrt{\kappa}}\right]$ we get
$\frac{J(r)}{r}\geq\frac{2}{\sqrt{\kappa}}\sin^{-1}\left(\sqrt{\frac{n+1}{2n}}\sin{\frac{\pi}{4}}\right)\frac{4\sqrt{\kappa}}{\pi}\geq\frac{8}{\pi}\sin^{-1}\left(\sqrt{\frac{1}{2}}\sin{\frac{\pi}{4}}\right)=\frac{8}{\pi}\frac{\pi}{6}=\frac{4}{3}.$
On the other hand, (3) implies that $\operatorname{diam}(A)\geq J(\operatorname{rad}(A))$ for $\operatorname{rad}(A)\in\left[0,\frac{\pi}{4\sqrt{\kappa}}\right]$. Hence, $\operatorname{diam}{A}\geq\frac{4}{3}\operatorname{rad}(A)$ as desired.
For Part (iii), we note from Theorem 3.3 that both the circumcenters $c(A)$ and $c(B)$ exist. The definition of $\operatorname{rad}(A)$ then implies that $A\subseteq\overline{\mathbb{B}(c(A),\operatorname{rad}{A})}$, the closed metric ball of radius $\operatorname{rad}(A)$ centered at $c(A)$. From the assumptions, we also have
$\operatorname{rad}(A)\leq\operatorname{diam}(A)<\Delta(X)\leq\rho(X).$
Therefore, $X^{\prime}\coloneqq\overline{\mathbb{B}(c(A),\operatorname{rad}{A})}$ is convex, hence a compact $\mathrm{CAT}(\kappa)$ space. Applying Theorem 3.3 for $B$ as a subset of $X^{\prime}$, we conclude that $c(B)\in X^{\prime}$. Therefore,
$d(c(B),c(A))<\operatorname{rad}(A)\leq\frac{3}{4}\operatorname{diam}(A).\qed$
Proposition A.1 (Path-connectedness).
Let $(X,d_{X})$ be path-connected and $(S,d_{S})$ an $(\varepsilon,R)$-approximation of $X$ for some $\varepsilon,R>0$. Then the geometric complex of $\mathcal{R}_{R+2\varepsilon}(S)$ is path-connected.
Proof.
Let $\mathscr{C}$ be an $(\varepsilon,R)$-correspondence between $S$ and $X$. Let $s,s^{\prime}\in S$ be arbitrary points. Then there exist points $x,x^{\prime}\in X$ such that $(s,x),(s^{\prime},x^{\prime})\in\mathscr{C}$. Since $X$ is assumed to be path-connected, so is $\mathcal{R}_{R}(X)$. As a result, there exists a sequence $\{x_{i}\}_{i=0}^{m+1}\subset X$ forming a path in $\mathcal{R}_{R}(X)$ joining $x$ and $x^{\prime}$. In other words, $x_{0}=x$, $x_{m+1}=x^{\prime}$, and $d_{X}(x_{i},x_{i+1})<R$ for $0\leq i\leq m$. There is also a corresponding sequence $\{s_{i}\}_{i=0}^{m+1}\subset S$ such that $s_{0}=s$, $s_{m+1}=s^{\prime}$, and $(s_{i},x_{i})\in\mathscr{C}$ for all $i$. We note that
$d_{S}(s_{i},s_{i+1})\leq d_{X}(x_{i},x_{i+1})+2\varepsilon<R+2\varepsilon.$
Thus, the sequence $\{s_{i}\}$ produces a path in $\mathcal{R}_{R+2\varepsilon}(S)$ joining $s$ and $s^{\prime}$. We conclude that the geometric complex of $\mathcal{R}_{R+2\varepsilon}(S)$ is path-connected. ∎